At a local restaurant, $287 in tips is to be shared between the server, bartender, and busboy. The server gets $21 more than three times the amount the busboy receives. The bartender gets $51 more than the amount the busboy receives. How much will the server receive?
Answer by rfer(16322)

bb=x=$43
bt=x+51=$94
s=3x+21=$150
---------------
x+x+51+3x+21=287
5x+72=287
5x=287-72
5x=215
x=215/5
x=43